It's a word problem, which I have no idea how to setup. I'm thinking it's a system of equations.
There were 44000 people at a ball game in Los Angeles. The day's receipts were $369,000. How many people paid $13.00 for a reserved seat and how many paid $6.00 for general admission?
Thank you
Chris

Found 2 solutions by JimboP1977, Alan3354:
Answer by JimboP1977(311)

369000 = 13x+6y (1)
44000 = x + y (2)
times (2) by 13 to give 572000 = 13x+13y (3)
(3) - (1)
203000 = 7y
y = 29000
therefore x must be 15000




Answer by Alan3354(69443)

g + r = 44000
6g + 13r = 369000
Can you finish it?